OpenStudy (anonymous):

A car has an acceleration of 2 m/s^2. What is its speed 3 seconds after it starts moving?

OpenStudy (anonymous):

i think this is derivative prob -are you in calc based class

OpenStudy (anonymous):

a = v/t 2 = v/3 v = 6